OpenStudy (anonymous):

Write √1694 in simplest form. Please help.

OpenStudy (lgbasallote):

\[\sqrt{1694} \implies \sqrt{2 \times 847} \implies \sqrt{2 \times 7 \times 121} \implies \sqrt{2 \times 7 \times 11 \times 11}\] do you see it?

OpenStudy (anonymous):

\[\sqrt{1694}=\sqrt{121*14}=11*\sqrt{14}\]

OpenStudy (compassionate):

41.15

OpenStudy (compassionate):

We all gave different answers. Lol. I'll give you the correct answer, one second.

OpenStudy (anonymous):

Please...

OpenStudy (compassionate):

√14 is the correct answer.

OpenStudy (lgbasallote):

lol they're all right

OpenStudy (compassionate):

I know they're all right. Haha. That's what's funny about it.

OpenStudy (lgbasallote):

what's funny is you were trying to make stanny believe they were wrong :p

OpenStudy (compassionate):

Oh, I didn't mean for it to come out like that. What I meant was I was going to narrow it down to what he was asking.

OpenStudy (anonymous):

lol I want to give the best answer to all of you

OpenStudy (lgbasallote):

lol you gave a NEW answer :p

OpenStudy (anonymous):

best response sorry

OpenStudy (compassionate):

Haha. Give it to whoever helped you the most.

OpenStudy (lgbasallote):

hmm you're not very new here are you @Stanny

OpenStudy (anonymous):

can i like give it to 2 people?

OpenStudy (compassionate):

No, you can't.

OpenStudy (anonymous):

very new, so fresh :D

OpenStudy (anonymous):

A sec please.

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=%E2%88%9A1694

OpenStudy (anonymous):

There you have the correct answer.